OpenStudy (anonymous):

The topic is the change in internal energy for systems by chemical reactions using the equation [DELTA]U = [DELTA]Hrxn + w "The change in internal energy is the change in heat of rxn + the work" This is the equation we learned in class . We were also taught that work = -[DELTA]nRT whether this is correct i dont know .. but this is what we were taught in class . So to calculate the change in internal energy we must subtract [DELTA]nRT from [DELTA]Hrxn , where R = the gas constant (8.314x10-3) and T will also be constant at 298K , atleast for this exam.

OpenStudy (aaronq):

if they told you that T=298K always, then yes.

OpenStudy (anonymous):

cont. (Im also assuming that atm will be constant at 1 , but im not sure if this even applies to this problem & was not even mentioned by my teacher) [DELTA]n would be the change in moles , as described by my teacher . Example Problem: N2(g) + 3H2(g) > 2NH3(g) --using hess law i am able to calculate [DELTA]Hrxn = -91.8kJ , now i plug this into my eqtn: [DELTA]U= -91.8kJ - [DELTA]nRT Now for [DELTA]nRT i think that [ mols of product (2) - mols of reactants (3+1=4) ] i can find [DELTA]n .. so [DELTA]nRT= (2-4)(8.314x10-3)(298) [DELTA]nRT= -4.95kJ so [DELTA]U = -98.1kJ + 4.95kJ [DELTA]U = -93.15KJ I dont know if this is the correct method , any help would be greatly appreciated !

OpenStudy (anonymous):

Thanks for the quick response , but you missed the second half of my question.

OpenStudy (aaronq):

It looks good to me

OpenStudy (anonymous):

are you sure of this, sir? my exam is tomorrow =(

OpenStudy (aaronq):

yes i am

OpenStudy (anonymous):

well thank you!

OpenStudy (aaronq):

you should have the actual answers in the back of your textbook, though. I suggest you look

OpenStudy (aaronq):

no problem

OpenStudy (anonymous):

not this type of problem. We were instructed to not read the book on this chapter, cause this guy wants us to do it his way. i guess. Horrible guy.